# On variants of Chebyshev’s conjecture

### v2

@@ 3.3. Proof of Theorem 2
First, we prove that the nonpositivity (1.10) for all sufficiently large $x>0$ implies the RH. Using the integral formula
@@
for $\Re(s)>1$. Combining this with (3.3) and denoting $2$ as $\log\exp(2)$, we obtain
$\int_{1}^{\infty}\left(-\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}\right)x^{-s+1/2}\,\frac{dx}{x}=-\frac{2(s-1)}{(s-1/2)^{2}}\frac{\zeta^{\prime}}{\zeta}(s)\quad\text{for}~\Re(s)>1.$ (3.6)
On the right-hand side, as stated in the proof of Theorem 1, $(\zeta^{\prime}/\zeta)(s)$ has no poles in the half-line $(1/2,\infty)$ except for the simple pole at $s=1$, but it is canceled out by s1s-1. Hence, it is shown that the nonpositivity (1.10) for sufficiently large $x>0$ implies the RH as in the proof of Theorem 1.
We prove the second half of the theorem. Applying (2.6) to $s=1/2$,
$2\sqrt{x}=\psi_{1/2}(x)+\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}+\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)-\sum_{k=1}^{\infty}\frac{x^{-2k-1/2}}{2k+1/2}$ (3.7)
for $x>1$. Substituting (3.7) into (3.1),
$\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-2\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}$ $=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)\log x+2\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}$ (3.8)
$\quad-\sum_{\rho}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}-\left(\frac{\zeta^{\prime}}{\zeta}\right)^{\prime}\left(\frac{1}{2}\right)+2\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)$
@@
By denoting $2$ as $\log\exp(2)$, the left-hand side of (3.8) is written as
@@
where the error term arises by replacing $\sideset{}{{}^{\prime}}{\sum}_{n\leq x}$ with $\sum_{n\leq x}$. Substituting this into (3.8), and then replacing $x$ with $xe^{2}$ and dividing by $\log x$, we have
$\sum_{n\leq xe^{2}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)$ $=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)+\frac{2}{\log x}\sum_{\rho}\frac{(xe^{2})^{\rho-1/2}}{\rho-1/2}$ (3.9)
@@
Suppose that equation (1.11) holds. Then, inequality (1.10) is valid for all sufficiently large $x>0$ by (3.2), and therefore, the RH holds from the first half of the theorem. From the RH, the third term on the right hand side of (3.9) is $O(1/\log x)$, and
@@
Hence, (1.11) implies that (1.12) holds. Conversely, assuming that the RH and (1.12) hold, (1.11) follows from (3.9). $\Box$
@@ 4.1. Proof of Theorem 3
$\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\Re(\chi(n))\log\frac{x}{n}$ (4.1)
has a constant sign for all $x\geq x_{0}$, and $L(\sigma)\not=0$ for $\sigma>\beta$ for some $1/2\leq\beta<1$. Then, $L(s)\not=0$ in the right-half plane $\Re(s)>\beta$. In particular, if (4.1) has a constant sign for all sufficiently large $x>0$, and $L(\sigma)\not=0$ for $\sigma>1/2$, then the GRH for $L(s)$ holds.
@@
$\lim_{x\to\infty}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{L^{\prime}}{L}\left(\frac{1}{2}\right)$ (4.2)
if and only if the GRH for $L(s,\chi)$ holds. If $\chi$ is primitive, then equation (4.2) implies that the sum (4.1) has a constant sign for all sufficiently large $x>0$.
@@
$\lim_{x\to\infty}\frac{1}{\log x}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{m}{2}$ (4.3)
if and only if the GRH for $L(s,\chi)$ holds. Equation (4.3) implies that the value of the sum (4.1) is negative for all sufficiently large $x>0$ if $m\geq 1$.
Remark. It is conjectured that $L(\sigma,\chi)\not=0$ for all $1/2\leq\sigma\leq 1$ and for all $\chi$. This conjecture is confirmed for primitive Dirichlet characters of modulus $q\leq 4\cdot 10^{5}$ by Platt [17], and for real odd Dirichlet characters of modulus $q\leq 3\cdot 10^{8}$ by Watkins [24].
@@
$f_{\chi}(x):=\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{n}.$ (4.4)
@@
$\int_{1}^{\infty}(-f_{\chi}(x))\,x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{(s-1/2)^{2}}\frac{L^{\prime}}{L}(s)\quad\text{for}~\Re(s)>1$ (4.5)
is obtained in the same way as (3.3) using the series expansion (2.9). On the right-hand side, there exists $1/2\leq\beta<1$ such that $L(\sigma)\not=0$ for any real number $\sigma>\beta$, because $L(s)$ is analytic in $\mathbb{C}$, and has no poles and zeros in the half-line $1,\infty)\subset\mathbb{R}$ by the Euler product (2.8) and Dirichlet’s theorem (2.7). Therefore, if $\chi$ is real, the first half of the theorem follows similarly to the proof of Theorem 1.
We consider the case where $\chi$ is not real. We replace $\chi$ by $\bar{\chi}$ in (4.5). Combining this with (4.5), we have
$\int_{1}^{\infty}\left(-\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\Re(\chi(n))\log\frac{x}{n}\right)$ $x^{-s+1/2}\,\frac{dx}{x}$ (4.6)
$=\frac{1}{(s-1/2)^{2}}\left(\frac{L^{\prime}}{L}(s,\chi)+\frac{L^{\prime}}{L}(s,\bar{\chi})\right).$
By $L(s,\bar{\chi})=\overline{L(\bar{s},\chi)}$, the assumption that $L(\sigma,\chi)\neq 0$ for $\sigma>\beta$ implies $L(\sigma,\bar{\chi})\neq 0$ for $\sigma>\beta$. This yields that the right-hand side of (4.6) has no poles on the interval $(\beta,\infty)\subset\mathbb{R}$. Thus, the assumption on (4.1) implies that the abscissa of convergence for the integral on the left-hand side of (4.6) is at most $\beta$ thanks to Proposition 1. This means that the right-hand side of (4.6) is analytic in $\Re(s)>\beta$. This forces $L(s,\chi)\neq 0$ in $\Re(s)>\beta$, since the residues of the logarithmic derivative at the zeros are positive integers and hence cannot cancel between $(L^{\prime}/L)(s,\chi)$ and $(L^{\prime}/L)(s,\bar{\chi})$.
@@
$\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{n^{s}}\log\frac{x}{n}$ $=-\frac{L^{\prime}}{L}(s)\log x-\sum_{\rho}\frac{x^{\rho-s}}{(\rho-s)^{2}}$ (4.7)
@@
for $x>1$, $s\not=\rho_{\chi},-2k-\kappa$ ($k\in\mathbb{Z}_{\geq 0}$). If $L(1/2)\not=0$, that is, if $m=0$, then substituting $s=1/2$ into (4.7) yields
$f_{\chi}(x)$ $=-\frac{L^{\prime}}{L}\left(\frac{1}{2}\right)\log x-\sum_{\rho}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}$ (4.8)
@@
$\lim_{s\to 1/2}$ $\left(-\frac{L^{\prime}}{L}(s)\log x-\frac{x^{-s+1/2}}{(-s+1/2)^{2}}-\left(\frac{L^{\prime}}{L}\right)^{\prime}(s)\right)$ (4.9)
@@
Therefore, by isolating the term of $\rho=1/2$ in the second sum on the right-hand side of (4.7), then taking the limit $s\to 1/2$, we obtain
$f_{\chi}(x)$ $=-\frac{m}{2}(\log x)^{2}-\lim_{s\to 1/2}\left[\frac{L^{\prime}}{L}(s,\chi)-\frac{m}{s-1/2}\right]\,\log x$ (4.10)
@@
$\quad-\sum_{\rho\not=1/2}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}-\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa^{\ast}-1/2}}{(2k+\kappa^{\ast}+1/2)^{2}},$
Assuming the GRH for $L(s)$, each term of the right-hand side (4.8) is bounded except for the first term, and each term of the right-hand side (4.10) is bounded except for the first and second terms. Hence, (4.2) and (4.3) holds.
Conversely, assuming (4.2) (resp. (4.3)), we have $f_{\chi}(x)\ll\log x$ (resp. $\ll(\log x)^{2}$) as $x\to\infty$. Therefore, the integral on the left-hand side of (4.5) converges uniformly on any compact subset in $\Re(s)>1/2$. Thus, it defines an analytic function in $\Re(s)>1/2$, and hence, $L(s)$ has no zeros (or poles) in $\Re(s)>1/2$, that is the GRH for $L(s)$.
Finally, we prove that the sum (4.1) has a constant sign for all sufficiently large $x>0$ assuming $L(1/2)\neq 0$, (4.2) and that $\chi$ is primitive. Taking the logarithmic derivative of (2.11) and substituting $s=1/2$, we obtain $\Re(\xi^{\prime}/\xi)(1/2,\chi)=0$. This gives
$\Re\frac{L^{\prime}}{L}\left(\frac{1}{2}\right)=\frac{1}{2}\left[\log\frac{\pi}{q}-\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{4}+\frac{\kappa(\chi)}{2}\right)\right]$ (4.11)
by (2.10). In particular, $\Re(L^{\prime}/L)(1/2)$ depends only on $q$ and the parity of $\chi$. By (4.11), the value $\Re(L^{\prime}/L)(1/2)$ vanishes only when $q>1$ takes the following values
@@
$\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{4}+\frac{\kappa(\chi)}{2}\right)=\left(\kappa(\chi)-\frac{1}{2}\right)\pi-3\log 2-C_{0}$ (4.12)
([15, p. 411]). Therefore, $\Re(L^{\prime}/L)(1/2)$ is a nonzero real number. Hence, by taking the real part of (4.2), it follows that (4.1) has a constant sign for all sufficiently large $x>0$. ∎
@@ 4.2. Proof of Theorem 4
$\sum_{p\leq x}\chi(p)\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}\leq 0$ (4.13)
@@
$\sum_{p\leq x}\chi(p)\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}=-\frac{\sqrt{x}}{4}(\log x)^{2}\left(1+O\left(\frac{1}{\log x}\right)\right)$ (4.14)
holds as $x\to\infty$. Therefore, (4.13) is valid for all sufficiently large $x>1$, and the left-hand side of (4.13) tends to -$\infty$ as $x\to\infty$.
@@
by using (2.1). For the third term on the right-hand side,
@@
by partial summation and (2.1). Therefore,
$f_{2}(x)=\frac{1}{4}(\log x)^{2}+O(\log x).$ (4.15)
@@
$f_{3}(x)=O(\log x)$ (4.16)
by the convergence of the sum. As a consequence, if we rewrite (4.5) as
$-\int_{1}^{\infty}f_{1}(x)\,x^{-s+1/2}\frac{dx}{x}$ $=\frac{1}{(s-1/2)^{2}}\frac{L^{\prime}}{L}(s)$ (4.17)
@@
Suppose that $f_{1}(x)$ has a constant sign for all sufficiently large $x>1$. Then, the assumption $L(\sigma,\chi)\not=0$ for $\sigma>\beta~(\geq 1/2)$ implies that the abscissa of convergence $\sigma_{c}$ for the integral on the left-hand side of (4.17) satisfies $\sigma_{c}\leq\beta$ by Proposition 1. Therefore, the integral on the left-hand side of (4.17) converges uniformly for any compact subset of $\Re(s)>\sigma_{c}$. This means that the right-hand side of (4.17) is analytic in $\Re(s)>\beta~(\geq 1/2)$. This forces $L(s,\chi)\neq 0$ in $\Re(s)>\beta~(\geq 1/2)$.
Conversely, suppose that the GRH for $L(s)$ holds. Then, $f_{\chi}(x)=O(\log x)$ by (4.8). On the other hand, $f_{\chi}(x)=f_{1}(z)+(1/4)(\log x)^{2}+O(\log x)$ by (4.15) and (4.16). Combining these two,
@@
This immediately implies (4.14). ∎

### v3

@@ 3.3. Proof of Theorem 2
First, we prove that the nonpositivity (10) for all sufficiently large $x>0$ implies the RH. Using the integral formula
@@
for $\Re(s)>1$. Combining this with (48) and denoting $2$ as $\log\exp(2)$, we obtain
$\int_{1}^{\infty}\left(-\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{ne^{2}}\right)x^{-s+1/2}\,\frac{dx}{x}=-\frac{2(s-1)}{(s-1/2)^{2}}\frac{\zeta^{\prime}}{\zeta}(s)\quad\text{for}~\Re(s)>1.$ (51)
On the right-hand side, as stated in the proof of Theorem 1, $(\zeta^{\prime}/\zeta)(s)$ has no poles in the half-line $(1/2,\infty)$ except for the simple pole at $s=1$, but it is canceled out by s1s-1. Hence, it is shown that the nonpositivity (10) for sufficiently large $x>0$ implies the RH as in the proof of Theorem 1.
We prove the second half of the theorem. Applying (36) to $s=1/2$,
$2\sqrt{x}=\psi_{1/2}(x)+\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}+\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)-\sum_{k=1}^{\infty}\frac{x^{-2k-1/2}}{2k+1/2}$ (52)
for $x>1$. Substituting (52) into (46),
$\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\log\frac{x}{n}-2\sideset{}{{}^{\prime}}{\sum}_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}$ $=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)\log x$ (53)
$\quad+2\sum_{\rho}\frac{x^{\rho-1/2}}{\rho-1/2}-\sum_{\rho}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}$
$\quad-\left(\frac{\zeta^{\prime}}{\zeta}\right)^{\prime}\left(\frac{1}{2}\right)+2\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)$
@@
By denoting $2$ as $\log\exp(2)$, the left-hand side of (53) is written as
@@
where the error term arises by replacing $\sideset{}{{}^{\prime}}{\sum}_{n\leq x}$ with $\sum_{n\leq x}$. Substituting this into (53), and then replacing $x$ with $xe^{2}$ and dividing by $\log x$, we have
$\sum_{n\leq xe^{2}}\frac{\Lambda(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)$ $=-\frac{\zeta^{\prime}}{\zeta}\left(\frac{1}{2}\right)+\frac{2}{\log x}\sum_{\rho}\frac{(xe^{2})^{\rho-1/2}}{\rho-1/2}$ (54)
@@
Suppose that equation (11) holds. Then, inequality (10) is valid for all sufficiently large $x>0$ by (47), and therefore, the RH holds from the first half of the theorem. From the RH, the third term on the right hand side of (54) is $O(1/\log x)$, and
@@
Hence, (11) implies that (12) holds. Conversely, assuming that the RH and (12) hold, (11) follows from (54). $\Box$
@@ 4.1. Proof of Theorem 3
$\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\Re(\chi(n))\log\frac{x}{n}$ (55)
has a constant sign for all $x\geq x_{0}$, and $L(\sigma)\not=0$ for $\sigma>\beta$ for some $1/2\leq\beta<1$. Then, $L(s)\not=0$ in the right-half plane $\Re(s)>\beta$. In particular, if (55) has a constant sign for all sufficiently large $x>0$, and $L(\sigma)\not=0$ for $\sigma>1/2$, then the GRH for $L(s)$ holds.
@@
$\lim_{x\to\infty}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{L^{\prime}}{L}\left(\frac{1}{2}\right)$ (56)
if and only if the GRH for $L(s)$ holds. If $\chi$ is primitive, then equation (56) implies that the sum (55) has a constant sign for all sufficiently large $x>0$.
@@
$\lim_{x\to\infty}\frac{1}{\log x}\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\left(1-\frac{\log n}{\log x}\right)=-\frac{m}{2}$ (57)
if and only if the GRH for $L(s)$ holds. Equation (57) implies that the value of the sum (55) is negative for all sufficiently large $x>0$ if $m\geq 1$.
Remark. It is conjectured that $L(\sigma,\chi)\neq 0$ for all $1/2\leq\sigma\leq 1$ and for all $\chi$. This conjecture has been confirmed for primitive Dirichlet characters of modulus $q\leq 4\cdot 10^{5}$ by Platt [17], and for real odd Dirichlet characters of modulus $q\leq 3\cdot 10^{8}$ by Watkins [24]. It has also been confirmed by Platt [17] that $L(1/2,\chi)\neq 0$ for primitive Dirichlet characters of modulus $q\leq 2\cdot 10^{6}$.
@@
$f_{\chi}(x):=\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{\sqrt{n}}\log\frac{x}{n}.$ (58)
@@
$\int_{1}^{\infty}(-f_{\chi}(x))\,x^{-s+1/2}\,\frac{dx}{x}=\frac{1}{(s-1/2)^{2}}\frac{L^{\prime}}{L}(s)\quad\text{for}~\Re(s)>1$ (59)
is obtained in the same way as (48) using the series expansion (39). On the right-hand side, there exists $1/2\leq\beta<1$ such that $L(\sigma)\not=0$ for any real number $\sigma>\beta$, because $L(s)$ is analytic in $\mathbb{C}$, and has no poles and zeros in the half-line $1,\infty)\subset\mathbb{R}$ by the Euler product (38) and Dirichlet’s theorem (37). Therefore, if $\chi$ is real, the first half of the theorem follows similarly to the proof of Theorem 1.
We consider the case where $\chi$ is not real. We replace $\chi$ by $\bar{\chi}$ in (59). Combining this with (59), we have
$\int_{1}^{\infty}\left(-\sum_{n\leq x}\frac{\Lambda(n)}{\sqrt{n}}\Re(\chi(n))\log\frac{x}{n}\right)$ $x^{-s+1/2}\,\frac{dx}{x}$ (60)
$=\frac{1}{2(s-1/2)^{2}}\left(\frac{L^{\prime}}{L}(s,\chi)+\frac{L^{\prime}}{L}(s,\bar{\chi})\right).$
By $L(s,\bar{\chi})=\overline{L(\bar{s},\chi)}$, the assumption that $L(\sigma,\chi)\neq 0$ for $\sigma>\beta$ implies $L(\sigma,\bar{\chi})\neq 0$ for $\sigma>\beta$. This yields that the right-hand side of (60) has no poles on the interval $(\beta,\infty)\subset\mathbb{R}$. Thus, the assumption on (55) implies that the abscissa of convergence for the integral on the left-hand side of (60) is at most $\beta$ thanks to Proposition 1. This means that the right-hand side of (60) is analytic in $\Re(s)>\beta$. This forces $L(s,\chi)\neq 0$ in $\Re(s)>\beta$, since the residues of the logarithmic derivative at the zeros are positive integers and hence cannot cancel between $(L^{\prime}/L)(s,\chi)$ and $(L^{\prime}/L)(s,\bar{\chi})$.
@@
$\sum_{n\leq x}\frac{\Lambda(n)\chi(n)}{n^{s}}\log\frac{x}{n}$ $=-\frac{L^{\prime}}{L}(s)\log x-\sum_{\rho}\frac{x^{\rho-s}}{(\rho-s)^{2}}$ (61)
@@
for $x>1$, $s\not=\rho_{\chi},-2k-\kappa^{\ast}$ ($k\in\mathbb{Z}_{\geq 0}$). If $L(1/2)\not=0$, that is, if $m=0$, then substituting $s=1/2$ into (61) yields
$f_{\chi}(x)$ $=-\frac{L^{\prime}}{L}\left(\frac{1}{2}\right)\log x-\sum_{\rho}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}$ (62)
@@
$\lim_{s\to 1/2}$ $\left(-\frac{L^{\prime}}{L}(s)\log x-\frac{mx^{-s+1/2}}{(-s+1/2)^{2}}-\left(\frac{L^{\prime}}{L}\right)^{\prime}(s)\right)$ (63)
@@
Therefore, by isolating the term of $\rho=1/2$ in the second sum on the right-hand side of (61), then taking the limit $s\to 1/2$, we obtain
$f_{\chi}(x)$ $=-\frac{m}{2}(\log x)^{2}-\lim_{s\to 1/2}\left[\frac{L^{\prime}}{L}(s,\chi)-\frac{m}{s-1/2}\right]\,\log x$ (64)
@@
$\quad-\sum_{\rho\not=1/2}\frac{x^{\rho-1/2}}{(\rho-1/2)^{2}}-\sum_{k=0}^{\infty}\frac{x^{-2k-\kappa^{\ast}-1/2}}{(2k+\kappa^{\ast}+1/2)^{2}}.$
Assuming the GRH for $L(s)$, each term of the right-hand side (62) is bounded except for the first term, and each term of the right-hand side (64) is bounded except for the first and second terms. Hence, (56) and (57) hold.
Conversely, assuming (56) (resp. (57)), we have $f_{\chi}(x)\ll\log x$ (resp. $\ll(\log x)^{2}$) as $x\to\infty$. Therefore, the integral on the left-hand side of (59) converges uniformly on any compact subset in $\Re(s)>1/2$. Thus, it defines an analytic function in $\Re(s)>1/2$, and hence, $L(s)$ has no zeros (or poles) in $\Re(s)>1/2$, that is the GRH for $L(s)$.
Finally, we prove that the sum (55) has a constant sign for all sufficiently large $x>0$ assuming $L(1/2)\neq 0$, (56) and that $\chi$ is primitive. Taking the logarithmic derivative of (41) and substituting $s=1/2$, we obtain $\Re(\xi^{\prime}/\xi)(1/2,\chi)=0$. This gives
$\Re\frac{L^{\prime}}{L}\left(\frac{1}{2}\right)=\frac{1}{2}\left[\log\frac{\pi}{q}-\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{4}+\frac{\kappa(\chi)}{2}\right)\right]$ (65)
by (40). In particular, $\Re(L^{\prime}/L)(1/2)$ depends only on $q$ and the parity of $\chi$. By (65), the value $\Re(L^{\prime}/L)(1/2)$ vanishes only when $q>1$ takes the following values
@@
$\frac{\Gamma^{\prime}}{\Gamma}\left(\frac{1}{4}+\frac{\kappa(\chi)}{2}\right)=\left(\kappa(\chi)-\frac{1}{2}\right)\pi-3\log 2-C_{0}$ (66)
([15, p. 411]). Therefore, $\Re(L^{\prime}/L)(1/2)$ is a nonzero real number. Hence, by taking the real part of (56), it follows that (55) has a constant sign for all sufficiently large $x>0$. ∎
@@ 4.2. Proof of Theorem 4
$\sum_{p\leq x}\chi(p)\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}\leq 0$ (67)
@@
$\sum_{p\leq x}\chi(p)\log p\cdot\sqrt{\frac{x}{p}}\,\log\frac{x}{p}=-\frac{\sqrt{x}}{4}(\log x)^{2}\left(1+O\left(\frac{1}{\log x}\right)\right)$ (68)
holds as $x\to\infty$. Therefore, (67) is valid for all sufficiently large $x>1$, and the left-hand side of (67) tends to -$\infty$ as $x\to\infty$.
@@
by using (31). For the third term on the right-hand side,
@@
by partial summation and (31). Therefore,
$f_{2}(x)=\frac{1}{4}(\log x)^{2}+O(\log x).$ (69)
@@
$f_{3}(x)=O(\log x)$ (70)
by the convergence of the sum. As a consequence, if we rewrite (59) as
$-\int_{1}^{\infty}f_{1}(x)\,x^{-s+1/2}\frac{dx}{x}$ $=\frac{1}{(s-1/2)^{2}}\frac{L^{\prime}}{L}(s)$ (71)
@@
Suppose that $f_{1}(x)$ has a constant sign for all sufficiently large $x>1$. Then, the assumption $L(\sigma,\chi)\not=0$ for $\sigma>\beta~(\geq 1/2)$ implies that the abscissa of convergence $\sigma_{c}$ for the integral on the left-hand side of (71) satisfies $\sigma_{c}\leq\beta$ by Proposition 1. Therefore, the integral on the left-hand side of (71) converges uniformly for any compact subset of $\Re(s)>\sigma_{c}$. This means that the right-hand side of (71) is analytic in $\Re(s)>\beta~(\geq 1/2)$. This forces $L(s,\chi)\neq 0$ in $\Re(s)>\beta~(\geq 1/2)$.
Conversely, suppose that the GRH for $L(s)$ holds. Then, $f_{\chi}(x)=O(\log x)$ by (62). On the other hand, $f_{\chi}(x)=f_{1}(z)+(1/4)(\log x)^{2}+O(\log x)$ by (69) and (70). Combining these two,
@@
This immediately implies (68). ∎